# supplement.pdf

### v2

@@ 4.1 Size
Table 2: Empirical sizes for Bernoulli exposures with a nonzero $\theta_{A}=c_{0}\textbf{1}_{q}$, where $c_{0}=0.5$ when $q=1$, and $c_{0}=0.3$ when $q=3$. $\Sigma_{0}=\Sigma_{AR}$ if not otherwise specified.
Sparsity of $\theta_{M}$ Case $q$ $n$ Zhou2020 Guo2022 Proposed
@@
$\beta_{A}=0$ 1 50 0.180 0.208 0.040
300 0.662 0.000 0.026
3 50 0.144 0.580 0.024
300 0.534 0.000 0.044
$\beta_{A}\neq 0$ 1 50 0.154 0.232 0.052
300 0.198 0.000 0.040
3 50 0.208 0.596 0.044
300 0.496 0.000 0.030
@@
1 50 0.122 0.152 0.054
300 0.688 0.082 0.040
3 50 0.108 0.336 0.026
300 0.746 0.100 0.050
Hard Sparsity ($\theta_{M}\neq 0$) $\beta_{A}=0$ 1 50 0.034 0.088 0.038
300 0.044 0.042 0.038
3 50 0.044 0.168 0.050
300 0.038 0.044 0.034
$\beta_{A}\neq 0$ 1 50 0.036 0.106 0.044
300 0.050 0.042 0.048
3 50 0.048 0.192 0.054
300 0.074 0.044 0.036
Capped-$\ell_{1}$ Sparsity ($\theta_{M}\neq 0$) $\beta_{A}=0$ 1 50 0.034 0.092 0.040
300 0.044 0.040 0.032
3 50 0.042 0.168 0.048
300 0.040 0.044 0.028
$\beta_{A}\neq 0$ 1 50 0.038 0.104 0.050
300 0.048 0.040 0.034
3 50 0.052 0.188 0.058
300 0.092 0.044 0.028
Decaying Coefficients ($\theta_{M}\neq 0$) $\beta_{A}=0$ 1 50 0.034 0.092 0.032
300 0.042 0.044 0.036
3 50 0.044 0.174 0.050
300 0.044 0.048 0.032
$\beta_{A}\neq 0$ 1 50 0.038 0.112 0.048
300 0.052 0.044 0.040
3 50 0.050 0.194 0.056
300 0.088 0.048 0.034
@@ 4.2 Power
In this subsection, we investigate the power behavior of the proposed test. The methods Zhou2020 and Guo2022 are not considered here since they are not always valid in our settings and thus a fair comparison in power is challenging. We adopt similar data generating processes as described in Section 4.1, but with different settings for $\beta_{A}$ and $\theta_{M}$. Specifically, given positive constants $c_{1},c_{2}\in\mathbb{R}$, we set the true coefficient vector $\theta_{M}$ according to the following three types of sparsity.
  • Hard sparsity: $\theta_{M,k}=0.3c_{2}k\mathbb{I}_{\left\{1\leq k\leq s\right\}}$, for $k=1,\ldots,p$;
  • Decaying coefficients: $\theta_{M,k}=c_{2}\left(0.3k\mathbb{I}_{\left\{1\leq k\leq s\right\}}+k^{-1.5}\mathbb{I}_{\left\{k\geq s+1\right\}}\right)$, for $k=1,\ldots,p$,
Also, set the true coefficient matrix $\beta_{A}$ with elements
@@
with $w_{k}\sim N(0,0.1^{2})$ for $k=1,\ldots,p$ and $j=1,\ldots,q$. Since both $c_{1}$ and $c_{2}$ are nonzero, the mediation effect $\gamma=\beta_{A}^{\top}\theta_{M}$ is nonzero and thus the alternative hypothesis of (3) is true. For simplicity we focus on the case of $q=1$ and consider two scenarios. In the first scenario, we fix $c_{2}=1$ (equivalent to fixing $\theta_{M}$) while vary $c_{1}$ (equivalent to varying $\beta_{A}$) with different values of $c_{1}\in\mathbb{R}$. In the second scenario, we fix $c_{1}=1$ (fix $\beta_{A}$) while vary $c_{2}$ (vary $\theta_{M}$) with different values of $c_{2}\in\mathbb{R}$. For each setting, the empirical power is computed as the proportion of rejections among the replications. Table 3 presents the simulation results of proposed test procedure when the exposures follow the Bernoulli distribution $\text{Bern}(0.5)$; the results for the Gaussian exposures are provided in Supplementary Material. As shown in the table, the empirical power increases as the signal (larger $c_{1}$ or $c_{2}$) or the sample size $n$ increases, numerically validating the results in Theorem 2.
@@
Case $n$ $c_{1}$ or $c_{2}$ Hard Sparsity Capped-$\ell_{1}$ Sparsity Decaying Coefficients
Fix $\theta_{M}$, Vary $\beta_{A}$ 50 1/8 0.080 0.084 0.082
1/4 0.158 0.160 0.152
1/2 0.420 0.422 0.422
1 0.824 0.824 0.822
300 1/8 0.150 0.148 0.140
1/4 0.526 0.526 0.520
1/2 0.926 0.920 0.920
1 0.998 0.998 0.998
Fix $\beta_{A}$, Vary $\theta_{M}$ 50 1/8 0.090 0.088 0.090
1/4 0.208 0.204 0.204
1/2 0.458 0.458 0.456
1 0.756 0.752 0.748
300 1/8 0.158 0.16 0.158
1/4 0.470 0.474 0.478
1/2 0.890 0.890 0.888
1 0.998 0.998 0.998
@@ 5 Data Application on LUAD
In recent decades, lung cancer has emerged as one of the most prevalent diseases worldwide (WHO, 2022; CDC, 2022; China, 2023), with smoking being a well-known key risk factor (Loeb et al., 1984; Shields, 2002; Vineis et al., 2004). Therefore, there is a pressing need to understand the underlying biological mechanisms that link smoking status and lung cancer development (Mok, 2011). Unraveling these mechanisms could provide invaluable insights for personalized prevention, diagnosis, and treatment strategies. DNA methylation, an epigenetic modification involving the addition of a methyl group to the DNA molecule, plays a crucial role in gene expression regulation and reflects various biological functions. Consequently, methylation markers are often considered potential mediators between exposures and health outcomes (Tobi et al., 2018; Kular et al., 2018; Fujii et al., 2022). In this section, our objective is to investigate how smoking exposure alters DNA methylation patterns, subsequently impacting an individual’s risk of developing lung cancer (Fasanelli et al., 2015; Zhang et al., 2016; Djordjilović et al., 2019). We utilize the data from The Cancer Genome Atlas (TCGA) project which provides a comprehensive characterization of multidimensional genomic alterations, including clinical and molecular data, across various cancer types (Tomczak et al., 2015). Specifically, we focus on the lung adenocarcinoma (LUAD) dataset from TCGA project using the R package cgdsr. To characterize the lung cancer development, we consider the forced expiratory volume in 1 second (FEV1), a commonly used measurement for assessing lung function in lung cancer studies (Young et al., 2007; Lange et al., 2015), as the outcome. The smoking status is encoded as a binary exposure $A$, where $A=1$ if the subject currently smokes or used to smoke. For each gene, the DNA methylation level measured by the Illumina Infinium HumanMethylation450 BeadChip is considered as a mediator. Instead of including all genes, it is more intuitive to investigate a set related genes each time. For this, we utilize the Gene Ontology (GO) terms that classify genes to groups according to the gene functions. Then, the DNA methylation levels of the genes in a GO term form a high-dimensional vector of mediators, and we apply the proposed test to each of the $7743$ GO terms from the Molecular Signatures Database (Subramanian et al., 2005; Liberzon et al., 2015), where the largest GO term contains $1781$ genes. In our analysis we also adjusts potential confounders including age, sex, and ethnicity, according to Section 2.3. After removing missing values, the number of subjects (the sample size $n$) in each GO term varies between $175$ and $211$. Before conducting the tests, we examine the structures of $\beta_{A}$ and $\theta_{M}$. For $\beta_{A}$, we have its least-squares estimator $\hat{\beta}_{A}=(\hat{\beta}_{A,j},j=1,\ldots,p)\in\mathbb{R}^{1\times p}$. To gain an insight into the sparsity of $\beta_{A}$, we compute the following surrogate
@@
where $\hat{\sigma}_{A,j}^{2}$ is the usual estimation of $\mathrm{Var}(\hat{\beta}_{A,j})$ in the linear regression model by regressing the $j$th mediator on the exposure and the potential confounders. Roughly speaking, $(\hat{\beta}_{A,j}-2\hat{\sigma}_{A,j},\hat{\beta}_{A,j}+2\hat{\sigma}_{A,j})$ forms an approximate confidence interval for $\beta_{A,j}$, and we consider $\beta_{A,j}$ to be nonzero if the interval does not contain 0. Therefore, $\hat{s}_{A}$ provides a rough estimate of the number of nonzero loadings of the vector $\beta_{A}$. For $\theta_{M}$, since we have its Lasso estimator $\hat{\beta}_{M}$, it is intuitive to estimate its sparsity by the number of nonzero loadings of $\hat{\beta}_{M}$, denoted by $\hat{s}_{M}$. Out of the 7743 datasets, we observe that,
  • 1338 datasets yield $\hat{s}_{A}=\hat{s}_{M}=0$, suggesting that a substantial proportion of the datasets likely fall into the case of $\beta_{A}=0$ and $\theta_{M}=0$;
@@
Consequently, it is crucial to adopt a test that remains valid in the above two cases, and the proposed test is the only option in the presence of high-dimensional mediators, to the best of our knowledge. We apply the test procedure proposed in Section 2.2, following the same implementation steps as in Section 4, to examine whether a candidate gene set associated with specific biological functions mediates the effect of smoking on lung function via the DNA methylation levels. By applying the proposed test to each of the aforementioned GO terms, we identify $169$ gene sets with significant mediation effects at the significance level of 5%, after a Bonferroni correction for multiple tests. Since many GO terms contain overlapping genes, it is not surprising to see this number of significant gene sets. Table 4 lists the top ten most significant gene sets. The majority of these gene sets are associated with immune response or cellular processes, suggesting that smoking may affect lung function through dysregulation of immune responses or disruption of cellular processes. These findings align with previous studies (Pfeifer et al., 2002; Stämpfli and Anderson, 2009), demonstrating the power of the proposed test for detecting mediation effect in the challenging setting of high-dimensional mediators.

### v3

@@ 4.1 Size
Table 2: Empirical sizes for Bernoulli exposures with a nonzero $\theta_{A}=c_{0}\textbf{1}_{q}$, where $c_{0}=0.5$ when $q=1$, and $c_{0}=0.3$ when $q=3$.
Sparsity for $\theta_{M}$ Case $q$ $n$ Zhou-00 Zhou-$1$ Guo-00 Guo-$1$ $\chi^{2}$-00 $\chi^{2}$-$1$ Bonf-00 Bonf-$1$
@@
$\beta_{A}=0$
($\Sigma_{0}=\Sigma_{AR}$)
1 50 0.180 0.164 0.208 0.208 0.056 0.046 0.056 0.046
300 0.662 0.652 0.000 0.000 0.034 0.032 0.034 0.032
3 50 0.144 0.122 0.580 0.580 0.036 0.018 0.040 0.022
300 0.534 0.512 0.000 0.000 0.040 0.032 0.048 0.042
$\beta_{A}\neq 0$-Sparse
($\Sigma_{0}=\Sigma_{AR}$)
1 50 0.154 0.152 0.232 0.232 0.066 0.052 0.066 0.052
300 0.198 0.196 0.000 0.000 0.038 0.036 0.038 0.036
3 50 0.208 0.192 0.596 0.594 0.040 0.038 0.064 0.042
300 0.496 0.496 0.000 0.000 0.030 0.026 0.028 0.028
$\beta_{A}\neq 0$-Dense
($\Sigma_{0}=\Sigma_{AR}$)
1 50 0.188 0.176 0.210 0.208 0.062 0.052 0.062 0.052
300 0.220 0.218 0.000 0.000 0.040 0.036 0.040 0.036
3 50 0.164 0.152 0.556 0.556 0.058 0.034 0.054 0.040
300 0.698 0.696 0.000 0.000 0.042 0.042 0.052 0.046
@@
1 50 0.122 0.110 0.152 0.152 0.082 0.068 0.082 0.068
300 0.688 0.686 0.082 0.078 0.042 0.040 0.042 0.040
3 50 0.108 0.098 0.336 0.336 0.072 0.048 0.070 0.036
300 0.746 0.742 0.100 0.098 0.062 0.058 0.058 0.050
$\beta_{A}\neq 0$-Dense
($\Sigma_{0}=\Sigma_{CS}$)
1 50 0.094 0.082 0.138 0.138 0.074 0.056 0.074 0.056
300 0.362 0.354 0.076 0.074 0.038 0.034 0.038 0.034
3 50 0.086 0.072 0.290 0.288 0.052 0.024 0.054 0.022
300 0.332 0.320 0.082 0.082 0.052 0.044 0.060 0.056
Hard ($\theta_{M}\neq 0$) $\beta_{A}=0$
($\Sigma_{0}=\Sigma_{AR}$)
1 50 0.034 0.034 0.088 0.088 0.038 0.034 0.038 0.034
300 0.044 0.044 0.042 0.042 0.042 0.042 0.042 0.042
3 50 0.044 0.042 0.168 0.166 0.054 0.050 0.056 0.054
300 0.038 0.038 0.044 0.044 0.034 0.034 0.036 0.034
$\beta_{A}\neq 0$-Sparse
($\Sigma_{0}=\Sigma_{AR}$)
1 50 0.036 0.036 0.106 0.106 0.048 0.044 0.048 0.044
300 0.050 0.050 0.042 0.042 0.050 0.046 0.050 0.046
3 50 0.048 0.048 0.192 0.192 0.046 0.046 0.062 0.056
300 0.074 0.074 0.044 0.044 0.044 0.042 0.038 0.036
$\beta_{A}\neq 0$-Dense
($\Sigma_{0}=\Sigma_{AR}$)
1 50 0.036 0.036 0.086 0.086 0.048 0.046 0.048 0.046
300 0.048 0.048 0.042 0.042 0.044 0.042 0.044 0.042
3 50 0.046 0.044 0.164 0.164 0.050 0.046 0.056 0.046
300 0.066 0.066 0.044 0.044 0.026 0.024 0.024 0.024
$\beta_{A}=0$
($\Sigma_{0}=\Sigma_{CS}$)
1 50 0.058 0.058 0.058 0.058 0.032 0.030 0.032 0.030
300 0.062 0.062 0.060 0.060 0.048 0.048 0.048 0.048
3 50 0.052 0.052 0.072 0.072 0.032 0.028 0.034 0.034
300 0.066 0.066 0.048 0.048 0.042 0.042 0.040 0.040
$\beta_{A}\neq 0$-Sparse
($\Sigma_{0}=\Sigma_{CS}$)
1 50 0.056 0.056 0.058 0.058 0.048 0.048 0.048 0.048
300 0.062 0.062 0.060 0.060 0.046 0.044 0.046 0.044
3 50 0.052 0.052 0.072 0.072 0.042 0.042 0.034 0.034
300 0.056 0.056 0.048 0.048 0.056 0.056 0.062 0.062
$\beta_{A}\neq 0$-Dense
($\Sigma_{0}=\Sigma_{CS}$)
1 50 0.058 0.058 0.058 0.058 0.036 0.034 0.036 0.034
300 0.060 0.060 0.060 0.060 0.050 0.050 0.050 0.050
3 50 0.050 0.050 0.074 0.074 0.036 0.034 0.038 0.038
300 0.072 0.072 0.048 0.048 0.046 0.046 0.046 0.044
@@ 4.2 Power
To examine the power behavior of the proposed test, we adopt data generating processes similar to those described in Section 4.1, but with different settings for $\beta_{A}$ and $\theta_{M}$. Specifically, given positive constants $c_{1},c_{2}\in\mathbb{R}$, we set the true coefficient vector $\theta_{M}$ with a hard-sparse structure: $\theta_{M,k}=0.3c_{2}k\mathbb{I}_{\left\{1\leq k\leq s\right\}}$ for $k=1,\ldots,p$. Other sparsity types are considered in Appendix E. Also, set the true coefficient matrix $\beta_{A}$ with elements
@@
with $w_{k}\sim N(0,0.1^{2})$ for $k=1,\ldots,p$ and $j=1,\ldots,q$. Since both $c_{1}$ and $c_{2}$ are nonzero, the mediation effect $\gamma=\beta_{A}^{\top}\theta_{M}$ is nonzero and thus the alternative hypothesis of (3) is true.
For simplicity we focus on the case of $q=1$ and consider two scenarios. In the first scenario, we fix $c_{2}=1$ (equivalent to fixing $\theta_{M}$) while vary $c_{1}$ (equivalent to varying $\beta_{A}$) with different values of $c_{1}\in\mathbb{R}$. In the second scenario, we fix $c_{1}=1$ (fix $\beta_{A}$) while vary $c_{2}$ (vary $\theta_{M}$) with different values of $c_{2}\in\mathbb{R}$. For each setting, the empirical power is computed as the proportion of rejections among the replications.
Table 3 presents the simulation results of proposed test Bond-$1$ when the exposures follow the Bernoulli distribution $\text{Bern}(0.5)$, in comparison with Zhou-$1$ and Guo-$1$; the results for the Gaussian exposures are provided in Appendix E. Note that $\chi^{2}$-$1$ is omitted as it is equivalent to Bond-$1$ when $q=1$. As shown in the table, the empirical power of Bond-$1$ increases as the signal (larger $c_{1}$ or $c_{2}$) or the sample size $n$ increases, numerically supporting the results in Theorem 3. Again, we observe that Zhou-$1$ and Guo-$1$ fail to control type-I errors in some cases of $c_{1}=0$ or $c_{2}=0$. In these cases, comparisons of power performance are not fair. However, in valid cases, Zhou-$1$ and Guo-$1$ demonstrate better power performance than the proposed test. Since the proposed test is designed to be valid across all null hypotheses, some loss of power is expected. It should also be noted that determining the validity of Zhou-$1$ and Guo-$1$ is already challenging, limiting the practical use of these methods. In contrast, the proposed method is more flexible and reliable.
@@
Case $n$ $c_{1}$ or $c_{2}$ Zhou-$1$ Guo-$1$ Bonf-$1$
Fix $\theta_{M}$, Vary $\beta_{A}$ 50 0 0.032 0.092 0.036
1/8 0.104 0.136 0.086
1/4 0.306 0.282 0.172
1/2 0.806 0.708 0.452
1 1 0.99 0.822
300 0 0.04 0.044 0.036
1/8 0.436 0.366 0.148
1/4 0.934 0.926 0.516
1/2 1 1 0.944
1 1 1 0.982
Fix $\beta_{A}$, Vary $\theta_{M}$ 50 0 0.12 0.246 0.04
1/8 0.726 0.306 0.104
1/4 0.92 0.528 0.252
1/2 0.944 0.814 0.492
1 0.946 0.926 0.764
300 0 0.114 0 0.038
1/8 0.988 0.552 0.162
1/4 1 1 0.468
1/2 1 1 0.932
1 1 1 0.994
@@ 5 Data Application on LUAD
In recent decades, lung cancer has emerged as one of the most prevalent diseases worldwide [33], with smoking being a well-known key risk factor [e.g., 34]. Therefore, there is a pressing need to understand the underlying biological mechanisms that link smoking status and lung cancer development [35]. Unraveling these mechanisms could provide invaluable insights for personalized prevention, diagnosis, and treatment strategies. DNA methylation, an epigenetic modification involving the addition of a methyl group to the DNA molecule, plays a crucial role in gene expression regulation and reflects various biological functions. Consequently, methylation markers are often considered potential mediators between exposures and health outcomes [e.g., 9]. In this section, our objective is to investigate how smoking exposure alters DNA methylation patterns, subsequently impacting an individual’s risk of developing lung cancer.
We utilize the data from The Cancer Genome Atlas (TCGA) project which provides a comprehensive characterization of multidimensional genomic alterations, including clinical and molecular data, across various cancer types [36]. Specifically, we focus on the lung adenocarcinoma (LUAD) dataset from TCGA project using the R package cgdsr. To characterize the lung cancer development, we consider the forced expiratory volume in 1 second (FEV1), a commonly used measurement for assessing lung function in lung cancer studies [e.g., 37], as the outcome. The smoking status is encoded as a binary exposure $A$, where $A=1$ if the subject currently smokes or used to smoke. For each gene, the DNA methylation level measured by the Illumina Infinium HumanMethylation450 BeadChip is considered as a mediator. Instead of including all genes, it is more intuitive to investigate a set related genes each time. For this, we utilize the Gene Ontology (GO) terms that classify genes to groups according to the gene functions. Then, the DNA methylation levels of the genes in a GO term form a high-dimensional vector of mediators, and we apply the proposed test to each of the $7743$ GO terms from the Molecular Signatures Database [38], where the largest GO term contains $1781$ genes. In our analysis we also include age, sex, and ethnicity as additional covariates, according to Remark 1 and Appendix B. After removing missing values, the number of subjects (the sample size $n$) in each GO term varies between $175$ and $211$.
Before conducting the tests, we examine the structures of $\beta_{A}$ and $\theta_{M}$. For $\beta_{A}$, we have its least-squares estimator $\hat{\beta}_{A}=(\hat{\beta}_{A,j},j=1,\ldots,p)\in\mathbb{R}^{1\times p}$. To gain an insight into the sparsity of $\beta_{A}$, we compute the following surrogate
@@
where $\hat{\sigma}_{A,j}^{2}$ is the usual estimation of $\mathrm{Var}(\hat{\beta}_{A,j})$ in the linear regression model by regressing the $j$th mediator on the exposure and the covariates. Roughly speaking, $(\hat{\beta}_{A,j}-2\hat{\sigma}_{A,j},\hat{\beta}_{A,j}+2\hat{\sigma}_{A,j})$ forms an approximate confidence interval for $\beta_{A,j}$, and we may consider $\beta_{A,j}$ to be nonzero if the interval does not contain 0. Therefore, $\hat{s}_{A}$ provides a rough estimate of the number of nonzero loadings of the vector $\beta_{A}$. For $\theta_{M}$, since we have its Lasso estimator $\hat{\theta}_{M}$, it is intuitive to estimate its sparsity by the number of nonzero loadings of $\hat{\theta}_{M}$, denoted by $\hat{s}_{M}$. Out of the 7743 datasets, we observe that,
  • 1338 datasets yield $\hat{s}_{A}=\hat{s}_{M}=0$, and for these datasets there is no strong evidence to rule out the case of $\theta_{M}=0$ and/or $\beta_{A}=0$;
@@
Consequently, it is crucial to adopt a test that remains valid in the above two cases, and the proposed test is the only option in the presence of high-dimensional mediators, to the best of our knowledge.
We apply the test procedure proposed in Section 2.2, following the same implementation steps as in Section 4, to examine whether a candidate gene set associated with specific biological functions mediates the effect of smoking on lung function via the DNA methylation levels. By applying the proposed test to each of the aforementioned GO terms, we identify $169$ gene sets with significant mediation effects at the significance level of 5%, after a Bonferroni correction for multiple tests. Since many GO terms contain overlapping genes, it is not surprising to see this number of significant gene sets. Table 4 lists the top ten most significant gene sets. The majority of these gene sets are associated with immune response or cellular processes, suggesting that smoking may affect lung function through dysregulation of immune responses or disruption of cellular processes. These findings align with previous studies [39], demonstrating the power of the proposed test for detecting mediation effect in the challenging setting of high-dimensional mediators.